# Retrieval-Enhanced Contrastive Vision-Text Models

### v1

@@ Abstract
Contrastive image-text models such as CLIP form the building blocks of many state-of-the-art systems. While they excel at recognizing common generic concepts, they still struggle on fine-grained entities which are rare, or even absent from the pre-training dataset. Hence, a key ingredient to their success has been the use of large-scale curated pre-training data aiming at expanding the set of concepts that they can memorize during the pre-training stage. In this work, we explore an alternative to encoding fine-grained knowledge directly into the model’s parameters: we instead train the model to retrieve this knowledge from an external memory. Specifically, we propose to equip existing vision-text models with the ability to refine their embedding with cross-modal retrieved information from a memory at inference time, which greatly improves their zero-shot predictions. Remarkably, we show that this can be done with a light-weight, single-layer, fusion transformer on top of a frozen CLIP. Our experiments validate that our retrieval-enhanced contrastive (RECO) training improves CLIP performance substantially on several challenging fine-grained tasks: for example +10.9 on Stanford Cars, +10.2 on CUB-2011 and +7.3 on the recent OVEN benchmark.
@@ 1 Introduction
In the recent years, we have witnessed a surge in the development of vision-language models that are highly adaptable to a broad spectrum of downstream tasks [27, 50, 65, 8, 55]. These models typically work by pre-training two parallel encoders using contrastive learning [46] on large-scale, carefully curated, image-text data [50]. These two-tower models learn to encode image and text into an aligned latent space which enables new appealing capabilities, such as zero-shot transfer to different downstream applications, e.g. image classification [50], image-text retrieval [48] or open-world recognition [44, 35]. Zero-shot transfer is a particularly effective approach to adapting to new tasks since the model is capable of handling new tasks without the need for fine-tuning on task-specific data or use of other domain adaptation protocols [63].
@@
Although these models have achieved state-of-the-art results across various generic vision-language benchmarks, our observation is that they tend to struggle on tasks requiring a more fine-grained understanding of visual or textual entities. Our hypothesis is that this disparity largely stems from the fact that during the pre-training phase, the contrastive loss compels the image representation to align with the representation of a short, sometimes noisy, text that neither exclusively nor comprehensively describes it. For example, we hypothesize that current vision-language models are good at associating images of cars with generic concepts such as “car”, “mechanics” or “road trip”, because these are common words paired with car images, but less at finegrained, instance-level, associations such as the specific brand, series or year of that car. This might therefore produce poor accuracy for zero-shot fine-grained car classification, where the task is to distinguish between different cars.
@@
Consequently, the current path taken by the research community has been to ever scale and curate the pre-training dataset in the hope of covering more and more, and cleaner image-text associations [50, 53, 1, 8]. However, one can wonder if this is a viable direction. An orthogonal effort has focused instead on memory or knowledge-based approaches [38, 24, 19, 26, 22, 37, 54]. These methods, instead of statically ingesting and memorizing all the world knowledge into model parameters, propose to rely on the access to an external source of knowledge. In this work, we follow their path and explore how existing vision-language models can leverage external knowledge for improved zero-shot fine-grained predictions. Also related to this goal, K-Lite [54] proposes to improve vision-text models by enhancing the text captions with more comprehensive text definitions retrieved from an external dictionary, i.e. WordNet [41] or Wiktionary [43]. Some caveats of this approach are (i) the retrieval is specific for the text-tower only, (ii) initial captions are augmented within their modality only, hence limiting the potential added-value of retrieval and (iii) the external knowledge base is limited, and may not contain all of the fine-grained entities of interest.
In this work, we explore retrieval-enhanced contrastive (RECO) vision-text models, i.e. a method to equip existing vision-text models with the ability to use relevant external knowledge in a cross-modal manner for improved zero-shot predictions. In particular, our method learns to refine original CLIP [50] text and visual embeddings with cross-modal items retrieved from a memory, namely a web-scale corpus of image-text pairs. For retrieval, we have observed that the image representation can be effectively utilized to identify relevant images closely resembling the query image, or the text representation can be used to identify relevant texts closely resembling the query text. By contrast, when crossing modalities, these representations are less successful in identifying suitable matches. We hence utilize the inherent strength of learned image and text representations within their respective modalities to retrieve relevant items and aid the alignment across modalities.
@@
Overall, as conceptually illustrated in Figure 1, we use an image representation as a query to identify the top-k most similar images and incorporate their associated text to create a multi-modal representation. In a parallel manner, given a text representation as a query, we find the top-k most similar texts and integrate their associated images to create a multi-modal representation. Through this process, we successfully transform the image and text representations into knowledge-enhanced multi-modal versions, which significantly enriches their representation power and simplifies their alignment in fine-grained zero-shot applications. This approach does not presuppose any downstream knowledge and produces a generic single model that can be used effectively across different downstream tasks. We validate our method through thorough design choice analyses. We show that our method improves over CLIP on 11 challenging fine-grained downstream tasks.
@@ 2 Related Work
While early works have shown the promise of representation learning from image-text paired data [67, 17, 28, 11], recent popular papers such as CLIP [50] and ALIGN [27] have truly unleashed the potential of contrastive image-text pre-training. This paradigm simply works with two parallel uni-modal encoders that learn to distinguish between aligned and non-aligned image-text pairs through a cross-modal contrastive objective [46, 42]. Appealing properties of these models are simplicity, scalability and great zero-shot performance [63]. As a result, vision-text contrastive models now form the basic building blocks of more powerful foundational models, such as CoCa [65], Flamingo [1], FLAVA [55], and PaLI [8] for example. In our work, we enhance the capabilities of the CLIP model [50], by adding a light-weight retrieval module. Nevertheless, our method is not specific to CLIP and can be applied to any vision-text model.
Knowledge-based vision-text models.
@@
Several works have focused on ways of improving upon different aspects of the contrastive vision-text models, such as their training objectives [15, 66, 12] or through scaling [9, 47]. Yet, only little exploration has been done on their combination with memory or knowledge-based techniques [14, 2, 37, 54]. REACT [37] retrieves image-text pairs from an external memory in order to build a training dataset specialized for a specific downstream task. Unlike REACT [37], our work does not require any pre-knowledge about the nature of the downstream task, and is hence applicable in a full zero-shot transfer. Another key difference is that our model can leverage items from the memory at inference time, while REACT uses retrieved items to automatically generate a training set to finetune their model. Closer to our work, K-LITE [54] learns vision-text models by leveraging external sources of knowledge (i.e. WordNet [41] or Wiktionary [43]) to complete captions with more descriptive content. Unlike our approach, the retrieved knowledge is uni-modal (e.g. they complement text with more text) and the external memory is not used for the image tower. Also using a knowledge-based approach but for image-only representation learning, NNCLR [14] finds the visual nearest-neighbor of each training image from a memory for contrastive learning. LGSimCLR [2] uses the language guidance to find most similar visual nearest-neighbor. Unlike our work, NNCLR and LGSimCLR only learn visual representations and use retrieval to enhance their supervision during training but not at inference.
Retrieval-based methods.
The main argument of the retrieval-based methods is that not all the world knowledge can be compiled into a model’s parameters. Thus, the model should also learn to rely on items retrieved from an external memory at inference. Retrieval-based methods have shown their promise in various NLP tasks [30, 21, 34, 60, 62, 4]. More recently, there is an increasing interest in the computer vision for retrieval-based methods as well [38, 24, 19, 26, 22, 37, 54]. Chen et al[7] and Blattmann et al[3] retrieve from an external memory for generative vision modeling. Long et al[38] and Iscen et al[25] propose retrieval-based methods for long-tailed classification. Hu et al[24] learn to retrieve from multiple external knowledge bases for visual question answering. Unlike these methods, our retrieval module does not need to be trained for a specific downstream task, and is capable of aggregating information for unseen tasks in a zero-shot manner.
@@ 3 Method
Preliminaries. We are given a pre-trained frozen dual-encoder vision-text model $f$, where $\mathbf{v}=f_{\text{image}}(I)$ is the embedding of image $I$, and $\mathbf{t}=f_{\text{text}}(T)$ is the embedding of text $T$. We say that these embeddings are uni-modal since they are obtained purely from a single modality, either image or text. We assume that image and text embedding spaces are already aligned, meaning that they have been trained to produce similar representations for matching image-text pairs and dissimilar representations for non-matching pairs [50, 27, 66, 46]. This alignment is usually obtained by minimizing the InfoNCE loss (or contrastive loss) [46] between embeddings of different modalities:
@@ 3.1 Retrieving cross-modal external knowledge
Cross-modal fusion. Our goal is to augment the text and visual original embeddings with cross-modal knowledge, not necessarily learned during the pre-training stage. For example, given the class name Yellow bellied flycatcher in a fine-grained bird classification problem such as CUB [59], we first look for captions in the memory that are semantically similar to the given class name. We then augment the class name representation with the visual representations of the retrieved similar captions, i.e. with what an Yellow bellied flycatcher looks like. Likewise, given a visual representation of a bird, we look for similar images in $\mathcal{M}$ and use their corresponding captions in the hope that some of them might contain useful information for our problem such as the species of that bird. Specifically, for a given text or image input, the retrieval module $\textsc{Knn}(.,\mathcal{M})$ returns items with the opposite modality than that of the input. We use the subscripts $v$ or $t$ to specify the modality of the retrieved embeddings. That is, $\textsc{Knn}_{t}(\mathbf{v},\mathcal{M})$ returns text embeddings from an image input and $\textsc{Knn}_{v}(\mathbf{t},\mathcal{M})$ returns image embeddings for text input.
Note that we also evaluate uni-modal fusion in our experiments, i.e. complementing visual representations with the retrieved visual knowledge and text representation with the retrieved captions. However, we find in practice that this variant leads to poorer performance than cross-modal fusion, as shown in Tab. 3. Intuitively, we hypothesize that this is because the signal brought by cross-modal fusion is richer due to the complementarity of the different modalities [25].
@@
We also evaluate cross-modal search but find that this leads to much poorer performance, especially in fine-grained problems, as shown in Tab. 3. A possible explanation is that the uni-modal search is an easier task, hence the retrieved elements are more relevant (because more similar) to the input. On the other hand, cross-modal search suffers from the pre-trained CLIP model’s lack of fine-grained alignment between the different modalities, resulting in noisier retrieval. Note that another advantage of uni- versus cross- modal search is that the latter requires the pre-trained image and text encoders to be already aligned while we can potentially let go of this hypothesis with uni-modal search. Finally, we illustrate the different scenarios for uni-modal versus cross-modal retrieval and fusion in Fig. 2 and quantitatively benchmark these scenarios in Tab. 3.
@@ 3.2 Learning how to fuse the retrieved knowledge
Transformer fusion. We model $\phi_{\text{image}}$ and $\phi_{\text{text}}$ as one-layer multi-head self-attention transformer encoders [58, 13]. Intuitively, this choice allows the original embedding to attend to all the retrieved elements in the fusion process. Note that while the fusion models for text and image encoders have identical architectures, they do not share parameters. In practice, the fusion module has a total of 3.16M parameters, which corresponds to only 2% of the total parameter count when using CLIP-B/32 as the backbone $f$. We have experimented with bigger fusion modules (see Appendix) but find that this light-weight solution works well in practice. We have also tried mean fusion of retrieved and original elements by simply averaging their embeddings but have found in practice that it performs poorly (see Tab. 3). Intuitively, the model needs to learn how to incorporate this new information, by, for example, learning how to omit or enhance some of the retrieved elements.
@@ 4 Experiments
In this section, we first detail our experimental protocol, then demonstrate the performance of RECO on several zero-shot benchmarks, and finally propose a thorough analysis of several design choices.
Table 1: Zero-shot transfer to image classification and retrieval. We report top-1 accuracy for classification and recall@1 for retrieval. We show the improvements obtained with RECO on top of CLIP-B/32 and CLIP-L/14: absolute performance gains are between brackets. For reference, we also include the performance of K-Lite [54] (another retrieval-augmented method) and other standard image-text foundation models (Flava[55], Align-base[27], LiT-L/16[66] or PaLI-17B [8]). We also report the total parameter count (“# par.”) of the different models (in Million).
@@
Method # par. Cars CUB Flowers Im1k Pl365 Dogs COCO Flickr COCO Flickr
CLIP-B/32 151 57.2 52.8 62.1 63.5 40.6 58.6 30.2 61.1 51.2 80.9
  + RECO 154 68.1(+10.9) 63.0(+10.2) 67.9(+5.8) 64.6(+1.1) 42.2(+1.6) 59.7(+1.1) 33.6(+3.4) 65.7(+4.6) 52.2(+1.1) 81.8(+0.9)
CLIP-L/14 428 75.6 61.7 75.6 75.5 42.0 72.7 35.2 68.6 57.2 87.5
  + RECO 435 82.8(+7.2) 73.4(+11.7) 79.5(+3.9) 76.1(+0.6) 43.6(+1.6) 73.9(+1.2) 38.7(+3.5) 72.6(+4.0) 58.0(+0.8) 88.5(+1.0)
@@
K-Lite[54] 151 10.0   – 78.6 52.3   –   –   –   –   –   –
Flava[55] 172   –   –   –   –   –   – 38.4 65.2 42.7 67.7
Align[27] 247 78.7 38.2 64.9 67.6 44.0 56.3 40.2 72.6 55.1 86.7
LiT-L[66] 638 24.8 61.7 74.1 75.8 42.6 70.8 31.1 57.6 48.5 77.7
PaLI[8] 17,000   –   –   – 72.1   –   –   –   –   –   –
@@
PaLI-17B [8] 17,000 4.4 1.2 1.9
CLIP-L/14 [50] 428 5.6 4.9 5.3
CLIP-L/14 [50] + RECO (Ours) 435 12.4 (+6.8) 12.7 (+7.8) 12.6 (+7.3)
@@
CLIP-L/14 Fusion [23] 880 33.6 4.8 8.4
PaLI-3B [8] 3,000 19.1 6.0 9.3
CLIP-L/14 CLIP2CLIP [23] 860 12.6 10.5 11.5
PaLI-17B [8] 17,000 28.3 11.2 16.1
4.1 Experimental setup
Training details. We train the fusion model on top of a frozen CLIP [50]. We use the CLIP-B/32 or the CLIP-L/14 versions. We train on Conceptual Captions 12M (“$\text{CC}_{\text{12M}}$”) [6], an image-text dataset containing about 10M pairs. We use a batch size of 4096, learning rate of $1e^{-3}$ decayed with a cosine schedule and weight decay of $1e^{-5}$. The temperature parameter is learned [50]. Training is done for 10 epochs, which lasts about 10 hours on a $44x$ 4 TPUv2 pod. For the memory, we use the subset of WebLI [8] containing 1B image-text pairs. We remove the near-duplicates of the test images from the memory. We have also explored using smaller but publicly available memory such as LAION-400M dataset [53] and show the results in Appendix.
[…]
Evaluation protocol. We evaluate in the zero-shot setting for all the considered benchmarks, meaning that no adaptation is done to the downstream task. As common in the literature [50, 27, 55, 66], we add prompts to the text of the downstream tasks, following [66]. We report the top-1 accuracy for all the tasks. We detail all evaluation protocols in Appendix.

### v2

@@ Abstract
Contrastive image-text models such as CLIP form the building blocks of many state-of-the-art systems. While they excel at recognizing common generic concepts, they still struggle on fine-grained entities which are rare, or even absent from the pre-training dataset. Hence, a key ingredient to their success has been the use of large-scale curated pre-training data aiming at expanding the set of concepts that they can memorize during the pre-training stage. In this work, we explore an alternative to encoding fine-grained knowledge directly into the model’s parameters: we instead train the model to retrieve this knowledge from an external memory. Specifically, we propose to equip existing vision-text models with the ability to refine their embedding with cross-modal retrieved information from a memory at inference time, which greatly improves their zero-shot predictions. Remarkably, we show that this can be done with a light-weight, single-layer, fusion transformer on top of a frozen CLIP. Our experiments validate that our retrieval-enhanced contrastive (RECO) training improves CLIP performance substantially on several challenging fine-grained tasks: for example +10.9 on Stanford Cars, +10.2 on CUB-2011 and +7.3 on the recent OVEN benchmark, where we even outperform the fine-tuned models on unseen classes.
@@ 1 Introduction
In the recent years, we have witnessed a surge in the development of vision-language models highly adaptable to a broad spectrum of downstream tasks (Jia et al., 2021; Radford et al., 2021; Yu et al., 2022; Chen et al., 2023; Singh et al., 2022). These models work by pre-training two parallel encoders using contrastive learning (van den Oord et al., 2018) on large-scale, carefully curated, image-text data (Radford et al., 2021). These two-tower models learn to encode images and texts into an aligned latent space which enables appealing capabilities such as zero-shot transfer to different downstream applications, e.g. image classification (Radford et al., 2021), image-text retrieval (Plummer et al., 2015) or open-world recognition (Minderer et al., 2022; Liang et al., 2022). Although these models have achieved state-of-the-art results across various generic vision-language benchmarks, we observe that they tend to struggle on tasks requiring a more fine-grained understanding of visual or textual entities. Our hypothesis is that this disparity stems from the fact that it is hard to align the image and text modalities. While every image is metaphorically valued at a thousand words, it is often paired with a short, sometimes noisy, text that neither exclusively nor comprehensively describes it. For example, current vision-language models are good at associating images of cars with generic concepts such as “car”, “mechanics” or “road trip”, because these are common words paired with car images, but less at finegrained, instance-level, associations such as the specific brand, series or year of that car. This might therefore produce poor accuracy for zero-shot fine-grained car classification.
@@
The current path taken by the research community has been to ever scale and curate the pre-training dataset in the hope of covering more and more image-text associations (Radford et al., 2021; Schuhmann et al., 2021; Alayrac et al., 2022; Chen et al., 2023). An orthogonal effort has focused instead on memory or knowledge-based approaches (Long et al., 2022; Hu et al., 2022; Gui et al., 2021; Izacard et al., 2022; Guu et al., 2020b; Liu et al., 2023; Shen et al., 2022). These methods, instead of statically ingesting and memorizing all the world knowledge into model parameters, propose to rely on the access to an external source of knowledge. For example, K-Lite (Shen et al., 2022) explores how to improve vision-text models by enhancing the text captions with more comprehensive text definitions retrieved from an external dictionary, i.e. WordNet (Meyer & Gurevych, 2012) or Wiktionary (Miller, 1998). One caveat that we identify in this approach is that initial captions are augmented within their modality only, hence limiting the potential added-value brought by the retrieved items.
To mitigate this issue, we put forth a retrieval-augmented approach that enhances the alignment between image and text representation. A critical observation of ours is that matching representations within the same modality is a significantly simpler task than matching representations across different modalities. To clarify, we observe that the image representation can be effectively utilized to identify images closely resembling the query image, or the text representation can be used to identify texts closely resembling the query text. However, when crossing modalities, these representations are less successful in identifying suitable matches, such as finding the text with the closest representation to a query image representation. We utilize the inherent strength of learned image and text representations within their respective modalities to aid the alignment across modalities. To improve their compatibility, we convert these unimodal representations into a multi-modal format, as conceptually illustrated in Fig. 1. Utilizing a web-scale corpus of image-text pairs for retrieval, we use image representation as a query to identify the top-k most similar images and incorporate the associated text to create a multi-modal representation. In a parallel manner, given a text representation as a query, we find the top-k most similar texts and integrate the associated images to create a multi-modal representation.
@@
Through this process, we successfully transform the image and text representations into multi-modal versions, which significantly simplifies their alignment. Our approach does not presuppose any downstream knowledge and produces a single generic model that can be used effectively across different tasks. We show that our method improves over original CLIP (Radford et al., 2021) or LiT (Zhai et al., 2022) models on 11 challenging fine-grained downstream tasks.
@@ 2 Related Work
Vision-text pre-training. While early works have shown the promise of representation learning from image-text paired data (Zhang et al., 2022; Gomez et al., 2017; Joulin et al., 2016; Desai & Johnson, 2021), recent popular papers such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) have truly unleashed the potential of contrastive image-text pre-training. This paradigm simply works with two parallel uni-modal encoders that learn to distinguish between aligned and non-aligned image-text pairs through a cross-modal contrastive objective (van den Oord et al., 2018; Miech et al., 2020). Appealing properties of these models are simplicity, scalability and great zero-shot performance (Xian et al., 2018). As a result, vision-text contrastive models now form the basic building blocks of more powerful foundational models, such as CoCa (Yu et al., 2022), Flamingo (Alayrac et al., 2022), FLAVA (Singh et al., 2022), and PaLI (Chen et al., 2023) for example. In our work, we enhance the capabilities of the CLIP model (Radford et al., 2021), by adding a light-weight retrieval module. Nevertheless, our method is not specific to CLIP and can be applied to any vision-text model.
@@
Knowledge-based vision-text models. Several works have focused on ways of improving upon different aspects of the contrastive vision-text models, such as their training objectives (Gao et al., 2022; Zhai et al., 2022; Dong et al., 2022) or through scaling (Cherti et al., 2022; Pham et al., 2021). Yet, only little exploration has been done on their combination with memory or knowledge-based techniques (Dwibedi et al., 2021; Banani et al., 2023; Liu et al., 2023; Shen et al., 2022; Fan et al., 2023). REACT (Liu et al., 2023) retrieves image-text pairs from an external memory in order to build a training dataset specialized for a specific downstream task. Unlike REACT (Liu et al., 2023), our work does not require any pre-knowledge about the nature of the downstream task, and is hence applicable in a full zero-shot transfer. Another key difference is that our model can leverage items from the memory at inference time, while REACT uses retrieved items to automatically generate a training set to finetune their model. Closer to our work, K-LITE (Shen et al., 2022) learns vision-text models by leveraging external sources of knowledge (i.e. WordNet (Meyer & Gurevych, 2012) or Wiktionary (Miller, 1998)) to complete captions with more descriptive content. Unlike our approach, the retrieved knowledge is uni-modal (e.g. they complement text with more text) and the external memory is not used for the image tower. Also using a knowledge-based approach but for image-only representation learning, NNCLR (Dwibedi et al., 2021) finds the visual nearest-neighbor of each training image from a memory for contrastive learning. LGSimCLR (Banani et al., 2023) uses the language guidance to find most similar visual nearest-neighbor. Unlike our work, NNCLR and LGSimCLR only learn visual representations and use retrieval to enhance their supervision during training but not at inference.
Retrieval-based methods. The main argument of the retrieval-based methods is that not all the world knowledge can be compiled into a model’s parameters. Thus, the model should also learn to rely on items retrieved from an external memory at inference. Retrieval-based methods have shown their promise in various NLP tasks (Khandelwal et al., 2020; Guu et al., 2020a; Lewis et al., 2020; Wang et al., 2022; Wu et al., 2022; Borgeaud et al., 2022). More recently, there is an increasing interest in the computer vision for retrieval-based methods as well (Blattmann et al., 2022; Fürst et al., 2022; Chen et al., 2022; Long et al., 2022; Hu et al., 2022; Gui et al., 2021; Izacard et al., 2022; Guu et al., 2020b; Liu et al., 2023; Shen et al., 2022; Iscen et al., 2023). Of particular interest, SuS-X (Udandarao et al., 2023) shows that by either retrieving similar samples to the query sample from a large data-bank like LAION can improve zero-shot classification performance of CLIP. Conceptually, SuS-X falls under the “Cross-modal search and cross-modal fusion” variant explored in this paper (see second scenario in Fig. 2). RA-CLIP (Xie et al., 2023) enriches the CLIP visual representation by retrieved image and text. However, their attempt to enrich the text representation degrades the performance, whereas we show that the retrieved data can also help produce better text representations.
@@ 3 Method
Preliminaries. We are given a pre-trained frozen dual-encoder vision-text model $f$, where $\mathbf{v}=f_{\text{image}}(I)$ is the embedding of image $I$, and $\mathbf{t}=f_{\text{text}}(T)$ is the embedding of text $T$. We say that these embeddings are uni-modal since they are obtained purely from a single modality, either image or text. We assume that image and text embedding spaces are already aligned, meaning that they have been trained to produce similar representations for matching image-text pairs and dissimilar representations for non-matching pairs (Radford et al., 2021; Jia et al., 2021; Zhai et al., 2022; van den Oord et al., 2018). This alignment is usually obtained by minimizing the InfoNCE loss (or contrastive loss) (van den Oord et al., 2018) between embeddings of different modalities:
@@ 3.1 Retrieving cross-modal external knowledge
Cross-modal fusion. Our goal is to augment the text and visual original embeddings with cross-modal knowledge, not necessarily learned during the pre-training stage. For example, given the class name Yellow bellied flycatcher in a fine-grained bird classification problem such as CUB (Wah et al., 2011), we first look for captions in the memory that are semantically similar to the given class name. We then augment the class name representation with the visual representations of the retrieved similar captions, i.e. with what an Yellow bellied flycatcher looks like. Likewise, given a visual representation of a bird, we look for similar images in $\mathcal{M}$ and use their corresponding captions in the hope that some of them might contain useful information for our problem such as the species of that bird. Specifically, for a given text or image input, the retrieval module $\textsc{Knn}(.,\mathcal{M})$ returns items with the opposite modality than that of the input. We use the subscripts $v$ or $t$ to specify the modality of the retrieved embeddings. That is, $\textsc{Knn}_{t}(\mathbf{v},\mathcal{M})$ returns text embeddings from an image input and $\textsc{Knn}_{v}(\mathbf{t},\mathcal{M})$ returns image embeddings for text input.
Note that we also evaluate uni-modal fusion in our experiments, i.e. complementing visual representations with the retrieved visual knowledge and text representation with the retrieved captions. However, we find in practice that this variant leads to poorer performance than cross-modal fusion, as shown in Tab. 3. Intuitively, we hypothesize that this is because the signal brought by cross-modal fusion is richer due to the complementarity of the different modalities (Iscen et al., 2023).
@@
We also evaluate cross-modal search but find that this leads to much poorer performance, especially in fine-grained problems, as shown in Tab. 3. An explanation is that the uni-modal search is an easier task, hence the retrieved elements are more relevant (because more similar) to the input. On the other hand, cross-modal search suffers from the pre-trained CLIP model’s lack of fine-grained alignment between the different modalities, resulting in noisier retrieval. Note that another advantage of uni- versus cross- modal search is that the latter requires the pre-trained image and text encoders to be already aligned while we can potentially let go of this hypothesis with uni-modal search.
@@ 3.2 Learning how to fuse the retrieved knowledge
Transformer fusion. We model $\phi_{\text{image}}$ and $\phi_{\text{text}}$ as one-layer multi-head self-attention transformer encoders (Vaswani et al., 2017; Dosovitskiy et al., 2021). Intuitively, this choice allows the original embedding to attend to all the retrieved elements in the fusion process. Note that while the fusion models for text and image encoders have identical architectures, they do not share parameters. In practice, the fusion module has a total of 3.16M parameters, which corresponds to only 2% of the total parameter count when using CLIP-B/32 as the backbone $f$. We have experimented with bigger fusion modules (see Appendix) but find that this light-weight solution works well in practice. We have also tried mean fusion of retrieved and original elements by simply averaging their embeddings but have found in practice that it performs poorly (see Tab. 3). Intuitively, the model needs to learn how to incorporate this new information, by, for example, learning how to omit or enhance some of the retrieved elements.
@@ 4 Experiments
4.1 Experimental setup
Training details. We train the fusion model on top of a frozen CLIP (-B/32 or -L/14 version) model (Radford et al., 2021). We also present a variant of RECO on top of a frozen LiT-L16L (Zhai et al., 2022) model. We train on Conceptual Captions 12M (“$\text{CC}_{\text{12M}}$”) (Changpinyo et al., 2021), an image-text dataset containing about 10M pairs. We use a batch size of 4096, learning rate of $1e^{-3}$ decayed with a cosine schedule and weight decay of $1e^{-5}$. The temperature parameter is learned (Radford et al., 2021). Training is done for 10 epochs, which lasts about 10 hours on a $44x$ 4 TPUv2 pod. For the memory, we use the subset of WebLI (Chen et al., 2023) containing 1B image-text pairs. We remove the near-duplicates of the test images from the memory. We have also explored using smaller but publicly available memory such as LAION-400M dataset (Schuhmann et al., 2021) and show the results in Appendix.
@@ 4.1 Experimental setup
Evaluation protocol. We evaluate in the zero-shot setting for all the considered benchmarks, meaning that no adaptation is done to the downstream task. As common in the literature (Radford et al., 2021; Jia et al., 2021; Singh et al., 2022; Zhai et al., 2022), we add prompts to the text of the downstream tasks, following (Zhai et al., 2022). All evaluation protocols are in Appendix.
Table 1: Zero-shot transfer to image classification. We report top-1 accuracy for classification. We show the improvements obtained with RECO on top of CLIP-R50, CLIP-B/32, CLIP-L/14 and LiT-L16L: absolute performance gains are between brackets. For reference, we also include the performance of K-Lite (Shen et al., 2022) and RA-CLIP (Xie et al., 2023) (other retrieval-augmented methods) and other image-text models (Align-base (Jia et al., 2021) and PaLI-17B (Chen et al., 2023)). We also report the total parameter count (“# par.”) of the different models (in Million).
Method # par. Cars CUB Flowers Im1k Pl365 Dogs
CLIP-R-50 102 38.6 52.0 47.2 59.2 53.1 60.6
  + RECO 114 39.8(+1.2) 62.8(+10.8) 56.2(+9.0) 59.4(+0.2) 54.0(+0.9) 64.4(+3.8)
CLIP-B/32 151 57.2 52.8 62.1 63.5 40.6 58.6
  + RECO 154 68.1(+10.9) 63.0(+10.2) 67.9(+5.8) 64.6(+1.1) 42.2(+1.6) 59.7(+1.1)
CLIP-L/14 428 75.6 61.7 75.6 75.5 42.0 72.7
  + RECO 435 82.8(+7.2) 73.4(+11.7) 79.5(+3.9) 76.1(+0.6) 43.6(+1.6) 73.9(+1.2)
LiT-L16L 638 90.5 54.5 77.4 80.2 45.2 75.7
  + RECO 652 90.8(+0.3) 74.8(+20.3) 84.1(+6.7) 80.9(+0.7) 45.4(+0.2) 81.3(+5.8)
@@
K-Lite 151 10.0   – 78.6 52.3   –   –
RA-CLIP 151   –   –   – 53.5   – 26.1
Align 247 78.7 38.2 64.9 67.6 44.0 56.3
PaLI 17,000   –   –   – 72.1   –   –
@@
PaLI-17B (Chen et al., 2023) 17,000 4.4 1.2 1.9
CLIP-L/14 (Radford et al., 2021) 428 5.6 4.9 5.3
CLIP-L/14 (Radford et al., 2021) + RECO (Ours) 435 11.5 (+5.9) 13.3 (+8.4) 12.3 (+7.0)
@@
CLIP-L/14 Fusion (Hu et al., 2023) 880 33.6 4.8 8.4
PaLI-3B (Chen et al., 2023) 3,000 19.1 6.0 9.3
CLIP-L/14 CLIP2CLIP (Hu et al., 2023) 860 12.6 10.5 11.5
PaLI-17B (Chen et al., 2023) 17,000 28.3 11.2 16.1